Actions
Run readiness discussion¶
To set the scene :
Today is the 6th Feb.
Saturday 11th : Expert led run
Wednesday week (15th) : Start data-taking
- Colin's Shutdown List : Hot off the press : Winter-Shutdown-_2716Tasks17-02-06.docx
- Hall Infrastructure
- Water : We think it's OK
- Power : Fine
- PPS : Fine
- Compressed air : Need to check with John on this issue. Critical to check.
- Target and Beamline
- Target and Target control PC : Target is fine - at risk until the PC is ready.
- Diffuser : Needs to be checked and contingency put in place
- Proton absorber : Henry gave it a good battering. Problem was a dodgy electrical connection.
- Conventional magnets : OK.
- Decay solenoid : Tucker claims it's ready to be powered. Get Mark to do it this week. Depends on compressed air -> need compressed air alarms in ALH.
- SS Magnets and Focus coil : Magnets are ready for action.
- Double check parameters for tags and make a new tag. Durga/Josef check tag creation script numbers and add currents for 2016-05-1 untrimmed unsettings.
- Detectors
- Trackers : HW mosty fine. Have ongoing issue concerning cooling performance. Some of the AFEs are running a bit hotter - regulation threshold has been increased.
Confusion in the last running when to do the tracker calibration daily <- MOM needs to remind shifters of this. - Hall probes
- EMR
- KL/CKOV
- TOFs
- Trackers : HW mosty fine. Have ongoing issue concerning cooling performance. Some of the AFEs are running a bit hotter - regulation threshold has been increased.
- Online systems
- Online computing : General environment OK. Network isolation will be tested on Saturday.
- DAQ : IOC for DAQmon got moved and buggered things. Hall probes are another dependency that the DAQ has. Needs to be fixed now.
- OnMon : works fine.
- OnRec : was running. Needs start/stop documentation.
- C&M inc. ALH
- Lots of spurious alarms which we need to look at.
- Will try to run from Pro - State machines need to be tested system-by-system.
- Stagegate for discussion on ALH is Friday morning @ 11 am.
- Saturday afternoon for expert led run.
- Operations
- Status of shifters and magnet monitors
- Spares list
- Online experts
- Shift checklist
- Magnet monitoring checklist
- AOB
- Environment spreadsheet should be implemented
- ISIS variables - check on Saturday
- Micestore space ALH - check on Saturday
- Update checklist
Updated by Boyd, Steven over 6 years ago ยท 8 revisions